What is the number missing in the sequence below?
5,13,23,37, (), 97

Answer by srinivas.g(540)

answer is 60
5= 2*2+1
13 = 3*3+4
23 =4*4+7
37 =5*5+12 (12=1+4+7)
60= 6*6+24 (24=1+4+7+12)
97 =7*7+48 (48=1+4+7+12+24)
so answer is 60